OpenStudy (anonymous):

The ratio of red marbles to blue marbles in a bag is 4:7. There are 16 red marbles. How many blue marbles are there?

OpenStudy (anonymous):

(Scroll to bottom if you just want the answer) The ratio given is \[\frac{ 4 }{ 7 }\] 4, the numerator, depicts the amount of red marbles. 7, the denominator, depicts the amounts of blue marbles. You now need to figure out how to get the numerator, 4, to 16. To do this you would multiply it by 4. Keep in mind that you must multiply the denominator by the same as the numerator! \[\frac{ 4*4 }{ 7*4 }\] The question is: If there are 16 red marbles, how many blue marbles are there? Since the remaining ratio is: \[\frac{ 16 }{ 28 }\] your answer would be 28. Final answer: 28 Blue marbles.

OpenStudy (anonymous):

Thank you.